OpenStudy (anonymous):

Could you please help me find the center, vertices, foci, eccentricity, retricemptotes, length of transverse axis and the length of the conjugate axis for \[\frac{ (x-4)^2 }{ 36 }-\frac{ (y-2)^2 }{ 9 }=1\]

OpenStudy (anonymous):

asymptotes*

OpenStudy (shamil98):

retricesymptotes LOL

OpenStudy (anonymous):

hahaha im tired and i am trying to finish 2 last questions for my study guide for the final in like 2 days o.o

OpenStudy (shamil98):

The center is found from the values (x-h)^2/a^2 - (y-k)^2/b^2 = 1 (h,k) --> center

OpenStudy (anonymous):

would the center be (4,2)?

OpenStudy (shamil98):

Eccentricity = c/a Yes that is the center

OpenStudy (anonymous):

Center: (4,2) eccentricity= c/a For the eccentricity you'll have to figure out what c is by plugging the values for a and be into the equation c^2=a^2-b^2

OpenStudy (shamil98):

Find C, from the equation a^2 = b^2 + c^2 solve for c. a^2 = 36, b ^2 = 9

OpenStudy (shamil98):

The length of the transverse axis is 2a, and the length of the conjugate axis is 2b.

OpenStudy (shamil98):

For the foci and vertices, I'm clueless atm about..

OpenStudy (anonymous):

okay for c^2 i got 27, so for the eccentricity it would be 27/36 which is .75?

OpenStudy (shamil98):

e = c /a c^2 = 27 c = 3√3

OpenStudy (anonymous):

for the lengths would it be 2(6) and 2(3)?

OpenStudy (shamil98):

a^2 = 36 a = 6 e = c/a e = 3√3/6 e = √3/2

OpenStudy (shamil98):

Yes. 12 , and 6. for the tranverse/conjugate axis

OpenStudy (anonymous):

where did you get 12 and 6 from?

OpenStudy (anonymous):

wouldn't it be 6 and 3?

OpenStudy (shamil98):

2a , and 2b i just multiplied them lol.

OpenStudy (anonymous):

OH hahaha

OpenStudy (anonymous):

thank you so much!!!

OpenStudy (anonymous):

do you know how to find the asymptotes?

OpenStudy (shamil98):

Nope.. I don't know how to find the foci/vertices either.

OpenStudy (anonymous):

its okay, thank you so much!! :)

OpenStudy (anonymous):

The foci are located at (4 (+/-) 3*sqrt(3), 2)